# PI-HMR: Towards Robust In-bed Temporal Human Shape Reconstruction with Contact Pressure Sensing

### v1

@@ 5.2 Ablations for PI-HMR
Table 3: Ablations for the K selection in TopK algorithm
@@
Table 4: Ablations for other components in MFF
Ablations for KD. We conduct experiments to evaluate cross-modal KD on our HMR pre-training. Tab. 5 shows that feature-based transfer plays a pivotal role in enhancing the performance, while Cliffs’ results might, to some extent, misguide the learning of HMR, due to domain gaps (Cliff’s encoder is pre-trained on ImageNet). When both supervisions coexist, HMR could learn the complete cognitive thought-chain of Cliff, leading to refinement in predictions.
GT Output-KD Feat-KD MPJPE PA-MPJPE
@@
Table 5: Ablations for cross-modal KD. GT, Output-KD, and Feat-KD represents supervision with GTs, CLiff’s outputs, and Cliff’s hidden feature maps, respectively.
@@ 5.3 Results for SMPLify-IB
2D MPJPE 3D Accel Limb height
Cliff 25.20 30.58 -
TIP 14.02 3.83 142.84
SMPLify-IB 9.65 6.77 66.68
Table 6: Qualitative results for SMPLify-IB, compared with the p-GTs in TIP, and CLIFF’s outputs. We calculate the 2D projection errors (in pixels), 3D acceleration (in $mm/s^{2}$), and the average height of limbs marked as stationary relative to the bed.
Tab. 6 provides the evaluation of p-GTs generated by SMPLify-IB. Besides the 2D projection errors and acceleration metrics, we introduces the static limb height as an objective assessment to our refinement in implausible limb lifts. Given the prevalence of limbs placed on other body parts within TIP, this metric can only serve as a rough estimate under limited self-penetration premise. We provide visual results in the Sup. Mat. to present our enhancements.
@@
We use SMPLify-XMC’s detection results as the GTs and conduct comparison experiments to evaluate our light-weight self-penetration detection algorithm in Tab. 7. The experiment run on the first group of the TIP dataset. For each batch with 128 images, we integrate both detection algorithm in our optimization routine, record the runtime for each iteration (1000 iterations for a batch) and calculate the accuracy, precision, and recall of the detection. Compared with SMPLify-XMC, our detection module achieves 53.9 times faster while maintaining a detection accuracy of 98.32%. We also implement a more lightweight version by domsampling the SMPL vertices into their 1/3 scale. The downsampled version further yields a more than tenfold increase in speed, accompanied by limited precision decrease.
@@ 6 Conclusion
In this work, we present a general framework for in-bed human shape estimation with pressure images, bridging from pseudo-label generation to algorithm design. For label generation, we present SMPLify-iB, a low-cost monocular optimization approach to generate SMPL p-GTs for in-bed scenes. By introducing gravity constraints and lightweight but efficient self-penetration detection module, we regenerate higher-quality SMPL labels for a public dataset TIP. For model design, we introduce PI-HMR, a pressure-based HPS network to predict in-bed motions from pressure sequence. By fusing pressure distribution and spatial priors, accompanying with KD and TTO exploration, PI-HMR outperforms previous methods. Results verify the feasibility of enhancing model’s performance by exploiting pressure’s nature.
@@ A Introduction
  • More qualitative results about SMPLify-IB, PI-HMR, and failure cases.
@@ C.2 Implementation details for PI-HMR
Before the aforesaid modules in the main body of our manuscript, PI-HMR also contains three different Transformer blocks for AttentionPooling, cross-attention for sampling features in MFF, and temporal consistency extraction. We will provide detailed designs of these Transformer layers. (1) For AttentionPooling, we use the same structure in CLIP [43]. (2) For cross-attention module in MFF, we apply a one-layer Transformer block as the attention module with one attention head and Dropout set as 0. (3) For the temporal encoder, we apply a two-layer Transformer block to extract the temporal consistency from the fusion feature sequence. In detail, each transformer layer contains a multi-head attention module with $N=8$ heads. These learned features are then fed into the feed-forward network with 512 hidden neurons. Dropout ($p=0.1$) and DropPath ($p_{d}=0.2$) are applied to avoid overfitting.
@@
$\mathcal{L}_{\text{SMPL}}=\omega_{s}^{\text{SMPL}}||\beta-\hat{\beta}||^{2}+\omega_{p}^{\text{SMPL}}||\theta-\hat{\theta}||^{2}+\omega_{t}^{\text{SMPL}}||t-\hat{t}||^{2})$
@@ C.5 Implementation details for Test-Time Optimization
We use the VQ-VAE to act as the only motion prior and supervision in our TTO routine. For terminological convenience, given a VQ-VAE $\mathbb{M}$ and PI-HMR initial predictions $\Theta^{n}_{0}=\{\theta^{n}_{1},...\theta^{n}_{T}\}$. For the $i_{th}$ iteration, we calculate the loss by Eq. 20 and update the $\Theta$ by stochastic gradient descent. The result of $i_{th}$ iteration will be input into $\mathbb{M}$ and optimized in the ${i+1}_{th}$ iteration:
$\mathcal{L}_{TTO}^{i}=\mathcal{L}_{m}(\Theta^{n}_{i},\Theta^{n}_{0})+\mathcal{L}_{m}(\Theta^{n}_{i},\mathbb{M}(\Theta^{n}_{i}))+\mathcal{L}_{sm}(\Theta^{n}_{i})$ (20)
@@
$\mathcal{L}_{m}(\Theta_{1},\Theta_{2})=\alpha\cdot$ $(\lambda_{\text{smpl}}^{TTO}||\Theta_{1},\Theta_{2}||^{2}$ (21)
$+\lambda_{3D}^{TTO}||\mathcal{J}(\Theta_{1})-\mathcal{J}(\Theta_{2})||^{2})$
@@ C.6 Implementation details for SOTA methods
D Visualization results
@@
D.1 Visualizations for Time Consumption of self-penetration algorithms
@@
D.2 Visualizations for gravity-based constraints.
@@
D.3 Failure cases for SMPLify-IB
@@ D.4 Failure cases for PI-HMR
D.5 More Qualitative Visualizations
@@
E Limitations and Future works

### v2

@@ 5.2 Ablations for PI-HMR
Table 3: Ablations for the K selection in TopK algorithm.
@@
Table 4: Ablations for other components in MFF.
Ablations for KD. We conduct experiments to evaluate cross-modal KD. Tab. 5 shows that feature-based transfer plays a pivotal role in enhancing the performance, while CLIFF’s results might, to some extent, misguide the learning of HMR, due to domain gaps (CLIFF’s encoder is pre-trained on ImageNet). When both supervisions coexist, HMR could learn the complete cognitive thought-chain of CLIFF, leading to refinement in predictions.
GT Output-KD Feat.-KD MPJPE PA-MPJPE
@@
Table 5: Ablations for cross-modal KD. GT, Output-KD, and Feat-KD represent supervision with GTs, CLIFF’s outputs, and CLIFF’s hidden feature maps, respectively.
@@ 5.3 Results for SMPLify-IB
2D MPJPE Limb height
CLIFF 25.20 -
TIP 14.02 142.84
SMPLify-IB 9.65 66.68
Table 6: Qualitative results for SMPLify-IB, compared with the p-GTs in TIP, and CLIFF’s outputs. We calculate the 2D projection errors (in pixels), and the average height of limbs marked as stationary relative to the bed.
Tab. 6 provides the evaluation of p-GTs generated by SMPLify-IB. Besides the 2D projection errors and acceleration metrics, we introduce the static limb height as an objective assessment of our refinement in implausible limb lifts. Given the prevalence of limbs placed on other body parts within TIP, this metric can only serve as a rough estimate under limited self-penetration premise. We provide visual results in the Sup. Mat. to present our enhancements.
@@
We use SMPLify-XMC’s detection results as the GTs and conduct comparison experiments to evaluate our light-weight self-penetration detection algorithm in Tab. 7. The experiment run on the first group of the TIP dataset. For each batch with 128 images, we integrate both detection algorithms in our optimization routine, record the runtime for each iteration (1000 iterations for a batch) and calculate the accuracy, precision, and recall of the detection. Compared with SMPLify-XMC, our detection module achieves 53.9 times faster while maintaining a detection accuracy of 98.32%. We also implement a more lightweight version by downsampling the SMPL vertices into their 1/3 scale. The downsampled version further yields a more than tenfold increase in speed, accompanied by limited precision decrease.
@@ 6 Conclusion
In this work, we present a general framework for in-bed human shape estimation with pressure images, bridging from pseudo-label generation to algorithm design. For label generation, we present SMPLify-IB, a low-cost monocular optimization approach to generate SMPL p-GTs for in-bed scenes. By introducing gravity constraints and a lightweight but efficient self-penetration detection module, we regenerate higher-quality SMPL labels for a public dataset TIP. For model design, we introduce PI-HMR, a pressure-based HPS network to predict in-bed motions from pressure sequences. By fusing pressure distribution and spatial priors, accompanied with KD and TTO exploration, PI-HMR outperforms previous methods. Results verify the feasibility of enhancing model’s performance by exploiting pressure’s nature.
@@
Figure 7: Our data flow includes three stages: (1) pre-train: knowledge distillation-based cross-modal pre-training; (2) train & evaluation: train the PI-HMR network with pressure sequences; (3) post-process: improve the estimates with the Test-Time Optimization strategy.
@@ A Introduction
  • More quantitative and qualitative results about SMPLify-IB, PI-HMR, and failure cases.
@@
The overall pipeline of our pressure-to-motion flow is shown in Fig. 7, and detailed architecture and implementation details will be elaborated below.
@@ C.2 Implementation details for PI-HMR
Before the aforesaid modules in the main body of our manuscript, PI-HMR also contains three different Transformer blocks for AttentionPooling, cross-attention for sampling features in MFF, and temporal consistency extraction. We will provide detailed designs of these Transformer layers. (1) For AttentionPooling, we use the same structure in CLIP [43]. (2) For the cross-attention module in MFF, we apply a one-layer Transformer block as the attention module with one attention head and Dropout set as 0. (3) For the temporal encoder, we apply a two-layer Transformer block to extract the temporal consistency from the fusion feature sequence. In detail, each transformer layer contains a multi-head attention module with $N=8$ heads. These learned features are then fed into the feed-forward network with 512 hidden neurons. Dropout ($p=0.1$) and DropPath ($p_{d}=0.2$) are applied to avoid overfitting.
@@
$\mathcal{L}_{\text{SMPL}}=\omega_{s}^{\text{SMPL}}||\beta-\hat{\beta}||^{2}+\omega_{p}^{\text{SMPL}}||\theta-\hat{\theta}||^{2}+\omega_{t}^{\text{SMPL}}||t-\hat{t}||^{2}$
@@ C.5 Implementation details for Test-Time Optimization
We use the VQ-VAE to act as the only motion prior and supervision in our TTO routine. For terminological convenience, given a VQ-VAE $\mathbb{M}$ and PI-HMR initial predictions $\Theta^{0}=\{\theta^{0}_{1},...\theta^{0}_{T}\}$. For the $i_{th}$ iteration, we calculate the loss by Eq. 20 and update the $\Theta$ by stochastic gradient descent. The result of $i_{th}$ iteration will be input into $\mathbb{M}$ and optimized in the ${i+1}_{th}$ iteration:
$\mathcal{L}_{TTO}^{i}=\alpha\mathcal{L}_{m}(\Theta^{i},\Theta^{0})+(1-\alpha)\mathcal{L}_{m}(\Theta^{i},\mathbb{M}(\Theta^{i}))+\mathcal{L}_{sm}(\Theta^{i})$ (20)
@@
$\mathcal{L}_{m}(\Theta_{1},\Theta_{2})=\lambda_{\text{smpl}}^{TTO}||\Theta_{1},\Theta_{2}||^{2}$ (21)
$+\lambda_{3D}^{TTO}||\mathcal{J}(\Theta_{1})-\mathcal{J}(\Theta_{2})||^{2}$
@@ C.6 Implementation details for SOTA methods
D More ablations
D.1 Discussion on TopK sampling
The sampling functions as a low-value filter, freeing the model’s attention from redundant, noisy backgrounds and focusing more on high-value regions. We provide a visualization in Fig. 10, where, with 128 points, the pressure image can retain the human’s outline while highlighting the core contact areas.
Figure 10: Visualization of TopK Sampling.
D.2 Comparisons with single-input models
For vision methods, single-image models usually exhibit lower MPJPE compared to temporal models (e.g. CLIFF vs PMCE). However, for pressure data, temporal models show superiority, likely due to their ability to leverage temporal context, mitigating information ambiguity. This implies the strength of temporal models in pressure data processing compared to single ones. For fair comparisons, we implemented a single-input-based PI-HMR, achieving a 62.01mm MPJPE (71.48mm for BodyMAP-WS), showing the efficacy of our architecture framework.
D.3 Results on the original TIP dataset
The results are shown in Tab. 9, which demonstrate a comparable magnitude of MPJPE reduction, proving the efficacy of PI-HMR.
Method TCMR PI-Mesh PI-HMR
MPJPE/ACC-ERR 67.9/14.6 79.2/18.2 68.38/5.24
Table 9: Quantitative results on the original TIP dataset.
D.4 Ablations of TTO.
We conducted ablations involving the selection of the balance weight $\alpha$ in Tab. 10 and the number of iterations in Tab. 11. We also explored integrating the pre-trained VQ-VAE into PI-Mesh during training (as it regresses the sequence rather than the mediate frame, making it suitable for VQ-VAE) and calculating the reconstruction loss. However, MPJPE drops limitedly (0.06mm). We will explore more potential methods (e.g. SPIN-like) in the future work.
$\alpha$ 0.1 0.3 0.5 0.7 0.9
MPJPE 56.94 55.93 55.50 55.43 55.67
Table 10: Ablations on balance weight $\alpha$.
iters 10 30 50 70 90
MPJPE 56.14 55.50 55.25 55.15 55.10
Table 11: Ablations on the number of iterations.
D.5 Generalization of SMPLify-IB on the SLP dataset.
We implemented SMPLify-IB on the SLP dataset. Results show the 2D MPJPE drops from 37.6 to 6.9 pixels compared to Cliff’s outputs. Fig. 11 shows our pros in alleviating depth ambiguity. Meanwhile, we observed limb distortions in the optimization results, which may stem from erroneous initial estimations (CLIFF exhibits notable domain adaptation issues in an in-bed scene). In the absence of temporal context, these mis-predictions could exacerbate the likelihood of unreasonable limb angles, underscoring the significance of temporal information in in-bed human shape annotations.
Figure 11: Visualizations of SMPLify-IB on SLP.
E Visualization results
@@
E.1 Visualizations for Time Consumption of self-penetration algorithms
@@
E.2 Visualizations for gravity-based constraints.
@@
E.3 Failure cases for SMPLify-IB
@@ E.4 Failure cases for PI-HMR
E.5 More Qualitative Visualizations
@@
F Limitations and Future works
@@
Acknowledgements: We thank the anonymous reviewers for their suggestions. This work is supported by the National Natural Science Foundation of China under Grant No. 62072420.